# On the parity of the Betti numbers of 3-manifolds with a parallel vector field

### v1

@@ Abstract.
The question of whether a closed, orientable manifold can admit a nonzero vector field that is parallel with respect to some Riemannian metric is a classical problem in Differential Geometry, first posed by S. S. Chern [chern1966geometry]. In this work, we provide a complete answer to Chern’s question in dimension three. Specifically, we show that a closed, orientable 3-manifold admits a nonzero parallel vector field with respect to some Riemannian metric if and only if it is a Kähler mapping torus. Furthermore, we prove that the Betti numbers of any such 3-manifold are necessarily odd. A full classification of these manifolds is also obtained.
@@ 1. Introduction
The purpose of this note is to resolves Chern’s question in dimension $3$. We proves that the only $3$-dimensional closed orientable manifolds that admit a nonvanishing parallel vector field with respect to a Riemannian metric are the Kähler mapping torus. As a consequence, we show that their Betti numbers are necessarily odd, and we provide a full classification.
@@ 3. Main Results
Let $M$ be a closed, orientable, 3-manifold. Then $M$ admits a nontrivial parallel vector field with respect to some Riemannian metric if and only if $M$ is a Kähler mapping torus.
@@
(Necessary part) $(\Rightarrow)$: Let $V$ be a parallel vector field on $M$ with respect to a Riemannian metric $g$, and let $\alpha:=i_{V}g$ denote the associated $1$-form. Denote by $D$ the Levi-Civita connection of $g$. Since $DV=0$, it follows that $D\alpha=0$, and consequently $d\alpha=0$. Furthermore, the function $g(V,V)$ is constant on $M$; we denote this constant by $g(V,V)=c>0$.
Normalize $V$ and $\alpha$ by setting
$\widetilde{V}:=\frac{V}{\sqrt{c}},\qquad\widetilde{\alpha}:=\frac{\alpha}{\sqrt{c}}.$
Then $\widetilde{V}$ and $\widetilde{\alpha}$ satisfy
$\widetilde{\alpha}(\widetilde{V})=1,\qquad D\widetilde{V}=0,\qquad\mathcal{L}_{\widetilde{V}}g=0,\qquad g(\widetilde{V},\widetilde{V})=1.$
Next, consider the $2$-form
$\Omega:=i_{\widetilde{V}}v_{g},$
where $v_{g}$ denotes the Riemannian volume form associated with $g$. A straightforward computation gives
$d\Omega=d(i_{\widetilde{V}}v_{g})=\mathcal{L}_{\widetilde{V}}v_{g}=0.$
Moreover, we observe that
$i_{\widetilde{V}}(\widetilde{\alpha}\wedge v_{g})=0,$
which in turn implies
$v_{g}=\widetilde{\alpha}\wedge\Omega.$
Hence, $\widetilde{V}$ is the Reeb vector field of the cosymplectic structure $(\widetilde{\alpha},\Omega)$. Since $\widetilde{V}$ is Killing, it follows from [bazzoni2015k] that $(M,\widetilde{\alpha},\Omega)$ is a $K$-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $M$ carries an almost contact metric structure $(\widetilde{\alpha},\phi,\widetilde{V},h)$ satisfying $L_{\widetilde{V}}h=0$. Furthermore, by [goldberg1969integrability, Proposition 3], this structure is normal ($N^{\phi}=0$), and hence $(M,\widetilde{\alpha},\widetilde{V},\phi,h)$ is a co-Kähler manifold. Finally, by [li2008topology, Theorem 2], $M$ is a Kähler mapping torus.
(Sufficient part) $(\Leftarrow)$: Consider a Kähler mapping torus $M=\Sigma_{\varphi}$. By [li2008topology, Theorem 2], $M$ admits a co-Kähler structure $(\phi,\xi,\eta,g)$ satisfying
@@
Moreover, the Nijenhuis tensor vanishes:
@@
As shown in [blair2010riemannian, blair1966theory, li2008topology, marrero1998new, cappelletti2013survey, olszak1981almost], every compact co-Kähler manifold satisfies
@@
where $\nabla^{LC}$ denotes the Levi-Civita connection associated with $g$. Since $\eta(X)=g(X,\xi)$, a straightforward calculation then shows that the Reeb vector field $\xi$ is parallel:
@@
This completes the first step of the proof. ∎
From Theorem 3.1, the only $3$-dimensional closed, orientable manifolds that admit a nonvanishing parallel vector field with respect to some Riemannian metric are the Kähler mapping torus. This completely answers Chern’s question in dimension $3$.
Yves Carrière [carriere1984flots] classified Riemannian flows on compact, orientable $3$-manifolds, and deduced that any such manifold admitting a nonvanishing Killing flow is a Seifert manifold. Now, consider a $3$-manifold $M$ endowed with a nonvanishing parallel vector field $V$ with respect to some Riemannian metric. Since $V$ is necessarily a Killing vector field, it follows from [carriere1984flots] that $M$ is a Seifert manifold. In what follows, we study the parity of the Betti numbers of $M$ and provide a complete classification of $3$-manifolds admitting a nonvanishing parallel vector field with respect to some Riemannian metric. This result may be viewed as a refinement of Carrière’s classification.
We can show the following result:
@@
A closed, orientable $3$-manifold $M$ admitting a nontrivial parallel vector field with respect to some Riemannian metric has odd Betti numbers, and is diffeomorphic to one of the following types:
@@
  2. (2)
    the three-dimensional torus $\mathbb{T}^{3}$;
  3. (3)
    one of the nontrivial torus bundles $\mathbb{T}^{2}_{A}$ over $\mathbb{S}^{1}$, where
@@
  4. (4)
    or a quotient of the form $(\mathbb{H}^{2}\times\mathbb{R})/\Gamma$, where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete subgroup acting freely.
@@
By Theorem 3.1, a closed, orientable manifold $M$ admitting a nontrivial parallel vector field with respect to some Riemannian metric is a Kähler mapping torus. According to [li2008topology, Theorem 2], such a manifold $M$ is co-Kähler manifold. Furthermore, by [Chinea1993TopologyOC, Theorem 11], the first Betti number of $M$ is odd. Applying Poincaré duality, it then follows that all Betti numbers of $M$ are odd. This concludes the first step of the proof.
@@
Since $M$ is a Kähler mapping torus, we have $M\simeq\Sigma_{\varphi}$, where $(\Sigma,J,h)$ is a Kähler manifold. Let $\mathrm{Isom}(\Sigma,h)$ denote the group of Hermitian isometries of $\Sigma$, and let $\mathrm{Isom}_{0}(\Sigma,h)$ denote its connected component of the identity. According to [bazzoni2014structure, Theorem 6.6], the diffeomorphism $\varphi$ is either isotopic to the identity or of finite order in the quotient group $\mathrm{Isom}(\Sigma,h)/\mathrm{Isom}_{0}(\Sigma,h)$; in particular, $\varphi$ is periodic. By [scott1983geometries, Theorem 5.4], $M$ is a Seifert bundle over a 2–orbifold $B$ with Euler number zero. Furthermore, as noted in [scott1983geometries], the three geometries $\mathbb{S}^{3}$, $\mathrm{Nil}$, and $\widetilde{\mathrm{SL}_{2}(\mathbb{R})}$ are excluded. Consequently, the classification of Seifert manifolds depends on the sign of the Euler characteristic $\chi(B)$. Case $\chi(B)>0$:
In this case, $M$ has $\mathbb{S}^{2}\times\mathbb{R}$ geometry. By [martelli2016introduction, Proposition 10.3.36], $M$ is diffeomorphic to one of $\mathbb{S}^{2}\times\mathbb{S}^{1}$, the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$, or $\bigl(S^{2};(p,q),(p,-q)\bigr)$. All manifolds of the type $\bigl(S^{2};(p,q),(p,-q)\bigr)$ are diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{1}$. Moreover, by [martelli2016introduction, Exercise 10.3.37], the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$ is diffeomorphic to $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$, which has first Betti number $b_{1}(M)=0$ and is therefore excluded. Hence, the only possibility is $M\simeq\mathbb{S}^{2}\times\mathbb{S}^{1}$. Case $\chi(B)=0$: Here, $M$ has $\mathbb{R}^{3}$ geometry, and by [martelli2016introduction, Theorem 12.3.1], $M$ admits a flat metric. From the classification results of Bieberbach, Hantzsche–Wendt, and Wolf [hantzsche1935dreidimensionale, Wolf, bieberbach1911bewegungsgruppen, bieberbach1912bewegungsgruppen], the first homology group $H_{1}(M,\mathbb{Z})$ is isomorphic to one of
$\mathbb{Z}^{3},\quad\mathbb{Z}\oplus\mathbb{Z}_{2}^{2},\quad\mathbb{Z}\oplus\mathbb{Z}_{3},\quad\mathbb{Z}\oplus\mathbb{Z}_{2},\quad\mathbb{Z},\quad\mathbb{Z}_{4}^{2}.$
Among these six flat manifolds, the only one with $b_{1}(M)=0$ is the Hantzsche–Wendt manifold, for which $H_{1}(M,\mathbb{Z})\cong\mathbb{Z}_{4}^{2}$; it is therefore excluded. All remaining manifolds have a positive first Betti number, and by [marrero1998new, Theorem 3.2], $M$ is diffeomorphic to one of
$M_{1}(1),\ M_{2}(1),\ M^{\prime}_{1}(1),\ M^{\prime}_{2}(1),\ \text{or }\mathbb{T}^{3},$
in the notation of [marrero1998new]. Moreover, as shown in the examples of [marrero1998new], the remaining manifolds are torus bundles, explicitly given by
$M_{1}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$,
$M_{2}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}$,
$M^{\prime}_{1}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}0&-1\\ 1&1\end{pmatrix}$,
$M^{\prime}_{2}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}-1&-1\\ 1&0\end{pmatrix}$.
Case $\chi(B)<0$: In this case, $M$ has geometry $\mathbb{H}^{2}\times\mathbb{R}$, and we have
$M\simeq(\mathbb{H}^{2}\times\mathbb{R})/\Gamma,$
where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete group acting freely [scott1983geometries].
This completes the proof.
@@
Let $M$ be a closed, orientable 3-manifold admitting a nontrivial parallel vector field $V$ with respect to some Riemannian metric $g$. Let $\alpha$ be the $1$-form metrically equivalent to $V$, and denote by $\mathcal{L}=\ker(\alpha)$ the corresponding foliation. Assume that the Ricci curvature of $g$ is non-negative on $\mathcal{L}\times\mathcal{L}$ and that $\mathcal{L}$ contains a compact leaf. Then the first Betti number $b_{1}(M)$ is either $1$ or $3$. Moreover, $b_{1}(M)=3$ if and only if $M$ is, up to a finite cover, diffeomorphic to the $3$-torus $\mathbb{T}^{3}$.
@@
Combining this with Theorem 3.2, we deduce that $b_{1}(M)$ is either $1$ or $3$. Finally, by [schliebner2015lorentzian, Theorem 1.1], the equality $b_{1}(M)=3$ holds if and only if $M$ is diffeomorphic to the $3$-torus $\mathbb{T}^{3}$. ∎

### v2

@@ Abstract.
The question of whether a closed, orientable manifold can admit a nontrivial vector field that is parallel with respect to some Riemannian metric is a classical problem in differential geometry, first posed by S. S. Chern [chern1966geometry]. In this work, we provide a complete answer to Chern’s question in dimension three. Specifically, we show that a closed, orientable 3-manifold admits a nontrivial parallel vector field with respect to some Riemannian metric if and only if it is a Kähler mapping torus. Furthermore, we prove that the Betti numbers of any such 3-manifold are necessarily odd. A full classification of these manifolds is also obtained. Similar results are established for compact orientable Lorentzian 3-manifolds admitting either a parallel timelike vector field or a parallel lightlike vector field.
@@ 1. Introduction
The purpose of this note is to resolve Chern’s question in dimension $3$. We prove that the only $3$-dimensional closed orientable manifolds that admit a nontrivial parallel vector field with respect to a Riemannian metric are Kähler mapping torus. As a consequence, we show that their Betti numbers are necessarily odd, and we provide a full classification. Analogous results are derived in the Lorentzian framework, where we consider $3$-manifolds endowed with either parallel timelike or lightlike vector fields.
@@ 3. Main Results
Let $M$ be a closed, orientable $3$-manifold. The following statements are equivalent:
  1. (1)
    $M$ admits a closed $1$-form $\alpha$ and a Killing vector field $Z$ such that
    $\alpha(Z)=\text{constant}\neq 0.$
  2. (2)
    $M$ is a Kähler mapping torus.
  3. (3)
    $M$ admits a nontrivial parallel vector field with respect to some Riemannian metric.
@@
We prove the equivalence by establishing the implications (1) $\Rightarrow$ (2), (2) $\Rightarrow$ (3), and (3) $\Rightarrow$ (1).
Proof that (1) $\Rightarrow$ (2). Assume that $M$ admits a closed $1$-form $\alpha$ and a Killing vector field $Z$ with respect to a Riemannian metric $g$, such that $\alpha(Z)$ is a nonzero constant. We may assume without loss of generality that $\alpha(Z)=1$. Let $v_{g}$ denote the Riemannian volume form associated with $g$, and define the $2$-form
$\Omega:=\iota_{Z}v_{g}.$
A straightforward computation gives
$d\Omega=d(\iota_{Z}v_{g})=\mathcal{L}_{Z}v_{g}=0,$
where the last equality follows from the fact that $Z$ is a Killing vector field. Moreover, we observe that
$\iota_{Z}(\alpha\wedge v_{g})=0,$
which implies
$v_{g}=\alpha\wedge\Omega.$
From the above, we deduce that $\alpha(Z)=1$ and $\iota_{Z}\Omega=0$. Hence, $Z$ is the Reeb vector field of the cosymplectic structure $(\alpha,\Omega)$. Since $Z$ is a Killing vector field with respect to $g$, it follows from [bazzoni2015k] that $(M,\alpha,\Omega,g)$ is a K-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $M$ carries an almost contact metric structure $(\eta,\phi,\xi,h)$ with $\eta=\alpha$ and $\xi=Z$, satisfying $\mathcal{L}_{Z}h=0$. Furthermore, by [goldberg1969integrability, Proposition 3], this structure is normal (i.e., the Nijenhuis tensor $N_{\phi}$ vanishes), and hence $(M,\alpha,Z,\phi,h)$ is a co-Kähler manifold. Finally, by [li2008topology, Theorem 2], $M$ is a Kähler mapping torus. This completes the proof of the first implication.
Proof that (2) $\Rightarrow$ (3). Suppose $M$ is a Kähler mapping torus. By [li2008topology, Theorem 2], $M$ admits a co-Kähler structure $(\phi,\xi,\eta,g)$ satisfying
@@
The metric $g$ is compatible with the almost contact structure $(\phi,\xi,\eta)$, and we have $\eta(X)=g(X,\xi)$. Moreover, the Nijenhuis tensor vanishes:
@@
As established in [blair2010riemannian, blair1966theory, li2008topology, marrero1998new, cappelletti2013survey, olszak1981almost], every compact co-Kähler manifold satisfies
@@
where $\nabla^{LC}$ denotes the Levi-Civita connection associated with $g$. Since $\eta(X)=g(X,\xi)$, a straightforward calculation shows that the Reeb vector field $\xi$ is parallel:
@@
This completes the proof of the second implication.
Proof that (3) $\Rightarrow$ (1). Assume that $M$ admits a nontrivial parallel vector field $Z$ with respect to a Riemannian metric $g$. Let $\beta:=\iota_{Z}g$ denote the $1$-form associated with $Z$, and let $D$ be the Levi-Civita connection of $g$. Since $DZ=0$, it follows that $D\beta=0$, and consequently $d\beta=0$. Furthermore, the function $g(Z,Z)=\beta(Z)$ is constant and positive on $M$, as $Z$ is nontrivial. The vector field $Z$ is a Killing vector field (since any parallel vector field is Killing) and $\beta(Z)$ is a positive constant. This completes the proof of the third implication and the theorem. ∎
From Theorem 3, the only $3$-dimensional closed, orientable manifolds that admit a nontrivial parallel vector field with respect to some Riemannian metric are the Kähler mapping torus. This completely answers Chern’s question in dimension $3$.
Yves Carrière [carriere1984flots] classified Riemannian flows on compact, orientable $3$-manifolds, and deduced that any such manifold admitting a nontrivial Killing flow is a Seifert manifold. Now, consider a $3$-manifold $M$ endowed with a nontrivial parallel vector field $Z$ with respect to some Riemannian metric. Since $Z$ is necessarily a Killing vector field, it follows from [carriere1984flots] that $M$ is a Seifert manifold. In what follows, we study the parity of the Betti numbers of $M$ and provide a complete classification of $3$-manifolds admitting a nontrivial parallel vector field with respect to some Riemannian metric. This result may be viewed as a refinement of Carrière’s classification.
@@
Let $M$ be a closed, orientable $3$-manifold admitting a nontrivial parallel vector field with respect to some Riemannian metric. Then:
@@
    2. (b)
      the $3$-torus $\mathbb{T}^{3}$;
    3. (c)
      a nontrivial torus bundle $\mathbb{T}^{2}_{A}$ over $\mathbb{S}^{1}$, where
@@
    4. (d)
      a quotient $(\mathbb{H}^{2}\times\mathbb{R})/\Gamma$, where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete subgroup acting freely.
@@
(1). By Theorem 3, $M$ is a Kähler mapping torus. According to [li2008topology, Theorem 2], such a manifold is co-Kähler. Furthermore, by [Chinea1993TopologyOC, Theorem 11], the first Betti number $b_{1}(M)$ is odd. Applying Poincaré duality, it follows that all Betti numbers $b_{k}(M)$ are odd.
@@
(2). Since $M$ is a Kähler mapping torus, we have $M\simeq\Sigma_{\varphi}$, where $(\Sigma,J,h)$ is a Kähler manifold. Let $\mathrm{Isom}(\Sigma,h)$ denote the group of Hermitian isometries of $\Sigma$, and $\mathrm{Isom}_{0}(\Sigma,h)$ its identity component. By [bazzoni2014structure, Theorem 6.6], the diffeomorphism $\varphi$ is either isotopic to the identity or of finite order in $\mathrm{Isom}(\Sigma,h)/\mathrm{Isom}_{0}(\Sigma,h)$; in particular, $\varphi$ is periodic. By [scott1983geometries, Theorem 5.4], $M$ is a Seifert fibered space over a 2-dimensional orbifold $B$ with Euler number zero. Moreover, as noted in [scott1983geometries], the geometries $\mathbb{S}^{3}$, $\mathrm{Nil}$, and $\widetilde{\mathrm{SL}_{2}(\mathbb{R})}$ are excluded. The classification now depends on the Euler characteristic $\chi(B)$ of the base orbifold. Case 1: $\chi(B)>0$. Here $M$ has $\mathbb{S}^{2}\times\mathbb{R}$ geometry. By [martelli2016introduction, Proposition 10.3.36], $M$ is diffeomorphic to one of: $\mathbb{S}^{2}\times\mathbb{S}^{1}$, the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$, or $\bigl(S^{2};(p,q),(p,-q)\bigr)$. All manifolds of type $\bigl(S^{2};(p,q),(p,-q)\bigr)$ are diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{1}$. By [martelli2016introduction, Exercise 10.3.37], the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$ is diffeomorphic to $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$, which has $b_{1}(M)=0$ and is therefore excluded. Hence, the only possibility is $M\simeq\mathbb{S}^{2}\times\mathbb{S}^{1}$. Case 2: $\chi(B)=0$. Here $M$ has Euclidean geometry. By [martelli2016introduction, Theorem 12.3.1], $M$ admits a flat metric. From the classification of closed flat 3-manifolds [hantzsche1935dreidimensionale, Wolf, bieberbach1911bewegungsgruppen, bieberbach1912bewegungsgruppen], the first homology group $H_{1}(M,\mathbb{Z})$ is isomorphic to one of: $\mathbb{Z}^{3}$, $\mathbb{Z}\oplus\mathbb{Z}_{2}^{2}$, $\mathbb{Z}\oplus\mathbb{Z}_{3}$, $\mathbb{Z}\oplus\mathbb{Z}_{2}$, $\mathbb{Z}$, or $\mathbb{Z}_{4}^{2}$. The Hantzsche–Wendt manifold (with $H_{1}(M,\mathbb{Z})\cong\mathbb{Z}_{4}^{2}$) has $b_{1}(M)=0$ and is excluded. All remaining manifolds have positive first Betti number, and by [marrero1998new, Theorem 3.2], $M$ is diffeomorphic to one of: $M_{1}(1)$, $M_{2}(1)$, $M^{\prime}_{1}(1)$, $M^{\prime}_{2}(1)$, or $\mathbb{T}^{3}$, in the notation of [marrero1998new]. Moreover, as shown in [marrero1998new], these are precisely the torus bundles: $M_{1}(1)\simeq\mathbb{T}^{2}_{A}$ with $A=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, $M_{2}(1)\simeq\mathbb{T}^{2}_{A}$ with $A=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}$, $M^{\prime}_{1}(1)\simeq\mathbb{T}^{2}_{A}$ with $A=\begin{pmatrix}0&-1\\ 1&1\end{pmatrix}$, and $M^{\prime}_{2}(1)\simeq\mathbb{T}^{2}_{A}$ with $A=\begin{pmatrix}-1&-1\\ 1&0\end{pmatrix}$. Case 3: $\chi(B)<0$. Here $M$ has $\mathbb{H}^{2}\times\mathbb{R}$ geometry, and thus $M\simeq(\mathbb{H}^{2}\times\mathbb{R})/\Gamma$, where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete subgroup acting freely [scott1983geometries].
This completes the proof of the theorem. ∎
The following corollaries are direct consequences of Theorem 3.2:
@@
Let $M$ be a closed, orientable 3-manifold admitting a nontrivial parallel vector field $Z$ with respect to some Riemannian metric $g$. Let $\alpha$ be the $1$-form metrically equivalent to $Z$, and denote by $\mathcal{L}=\ker(\alpha)$ the corresponding foliation. Assume that the Ricci curvature of $g$ is non-negative on $\mathcal{L}\times\mathcal{L}$ and that $\mathcal{L}$ contains a compact leaf. Then the first Betti number $b_{1}(M)$ is either $1$ or $3$. Moreover, $b_{1}(M)=3$ if and only if $M$ is, up to a finite cover, diffeomorphic to the $3$-torus $\mathbb{T}^{3}$. If $b_{1}(M)=1$ then the flow of $Z$ has either exactly two closed orbits or infinitely many closed orbits. In the first case $M$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{1}$, and in the second case $M$ is either a nontrivial $\mathbb{T}^{2}$-bundle over $\mathbb{S}^{1}$ or a quotient $(\mathbb{H}^{2}\times\mathbb{R})/\Gamma$, where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete subgroup acting freely.
@@
Combining this with Theorem 3.2, we deduce that $b_{1}(M)$ is either $1$ or $3$. Moreover, by [schliebner2015lorentzian, Theorem 1.1], the equality $b_{1}(M)=3$ holds if and only if $M$ is, up to a finite cover, diffeomorphic to the $3$-torus $\mathbb{T}^{3}$. Now assume that $b_{1}(M)=1$. We distinguish two cases according to whether the number of closed orbits of $Z$ is finite or infinite. In the finite case, as noted in the comments following [bazzoni2015k, Corollary 8.7], the number of closed Reeb orbits $N$ is given by $\dim H^{*}(M,\mathcal{F}_{Z})$, where $H^{*}(M,\mathcal{F}_{Z})$ is the basic cohomology of $M$ with respect to the foliation $\mathcal{F}_{Z}$ induced by $Z$. By Theorem 3, the manifold $M$ is K-cosymplectic. According to [bazzoni2015k, Theorem 7.6], when $b_{1}(M)=1$, the $\mathbb{S}^{1}$-action is automatically Hamiltonian (this fact is also noted on p. 31 of [bazzoni2015k]). For a 3-dimensional K-cosymplectic manifold with $b_{1}(M)=1$, [bazzoni2015k, Corollary 8.7] gives
$N=1+b_{1}(M)=2.$
From [carriere1984flots, Théorème 1], it follows that $M$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{1}$. The infinite case follows directly from Theorem 3.2. ∎
Corollary 3.2.
Let $(M,g)$ be a compact, orientable Lorentzian $3$-manifold. Then $M$ has odd Betti numbers and is diffeomorphic to one of the manifolds listed in Theorem 3.2 if either of the following holds:
  1. (1)
    $M$ admits a timelike vector field $Z$ which is parallel with respect to $g$; or
  2. (2)
    $M$ admits a closed $1$-form $\alpha$ and a lightlike vector field $Z$ with $\alpha(Z)=\mathrm{const}\neq 0$, where $Z$ is parallel with respect to $g$.
Proof.
(1). Suppose $M$ admits a timelike vector field $Z$ that is parallel with respect to the Lorentzian metric $g$. Normalize so that g(Z,Z)=1g(Z,Z)=-1, and let $\beta=\iota_{Z}g$ be the associated 1-form. Define a Riemannian metric $h$ by
$h=g+2\beta\otimes\beta.$
Then $h(Z,Z)=1$, and since $Z$ is parallel with respect to $g$, the metrics $g$ and $h$ share the same Levi–Civita connection. Hence, $Z$ is also parallel with respect to $h$. The conclusion now follows from Theorem 3.2.
(2). Suppose $M$ admits a closed 1-form $\alpha$ and a lightlike vector field $Z$ that is parallel with respect to the Lorentzian metric $g$, with $\alpha(Z)=\text{constant}\neq 0$. By [zeghib1996killing, Section 14], $Z$ is a Killing vector field with respect to some Riemannian metric. Theorem 3 then implies that $M$ admits a nontrivial parallel vector field with respect to some Riemannian metric, and the result follows from Theorem 3.2. ∎